OpenStudy (anonymous):

integral of (sin 3t + cos 3t)^2

OpenStudy (anonymous):

\[f(t) = (\sin (3) *t + \cos (3) *t)^{2}\]\[\int\limits f (t) dx = (\sin (3) *t + \cos (3) *t ^{2} 8 x\]

OpenStudy (loser66):

Manipulate the integrand a little bit: \((sin (3t) +cos (3t))^2= sin^2(3t) + 2 sin(3t) cos(3t) + cos^2(3t)=1+ sin(6t)\) Now, you can take integral term by term.

OpenStudy (anonymous):

Thank you